OpenStudy (anonymous):

Which set of side lengths defines a right triangle? A. 7, 9, 13 B. 5, 12, 14 C. 8, 15, 17 D. 9, 12, 14

TheSmartOne (thesmartone):

@Michele_Laino @nincompoop

TheSmartOne (thesmartone):

@Abhisar QH help :)

OpenStudy (anonymous):

thank you. :)

OpenStudy (anonymous):

simple rule a^2 + b^2 = c^2

OpenStudy (michele_laino):

the right set has to be a Pitagora's set

OpenStudy (michele_laino):

hint: we have: 17^2= 8^2+ 15^2

OpenStudy (anonymous):

You can check every set, is the condition valid for it?